concentrated phosphoric acid solution is 85.5% H3PO4 by mass and has a density of 1.69 g/mL at 25°C. What is the molarity of H3PO4? What is the mole fraction of urea, CO(NH2)2, in a solution prepared by dissolving 5.6 g of urea in 30.1 g of methanol, CH3OH? How will an understanding of this concept help you in your healthcare career?
Answer #1

1)

H3PO4 molar mass = 98 g / mole

mass % = 85.5 %

density = 1.69 g / mL

formula : relation between mass % and molarity

molarity = mass % x 10 x density / molar mass

             = 85.5 x 10 x 1.69 / 98

             = 14.7 M

molarity = 14.7 M

2)

moles of urea = mass / molar mass

                       = 5.6 / 60

                        = 0.0933

moles of methanol = 30.1 / 32

                             = 0.961

mole fraction of urea = moles of urea / total moles

                                   = 0.0933 / 0.0933 + 0.961

                                  = 0.0902

mole fraction of urea = 0.0902